Two charged pellets, each with a charge of 5C, are located a distance 15m apart. What is the electric force between the two pellets?
Answer #1

Here,

charge , Q = 5 C

distance , d = 15 m

for the electric force between two pellets

electric force = k * Q^2/d^2

electric force = 9 *10^9 * 5^2/15^2

electric force = 1 *10^9 N

the electric force between the pellets is 1 *109 N